OpenStudy (anonymous):

integral of x^2 cos(X^3) dx

OpenStudy (anonymous):

try \(u=x^3\)

OpenStudy (ash2326):

@ncorrea92 can you do it now?

OpenStudy (anonymous):

-1/3 sin (X^3) + c is that the answer?

OpenStudy (ash2326):

Oh just a silly mistake integral of cos x is sin x so \[\frac{\sin x^3}{3} +c\]

OpenStudy (ash2326):

Great work though:)